Given h(x) = 2x + 9 and f(x) = 1/6x +5 Evaluate the composite function h[f(6)]

Given

h(x) = 2x + 9
and
f(x) = 1/6x +5

Evaluate the composite function h[f(6)]
Answer #1

Answer:  h[f(6)]=21

Explanation:

h(x)=2x+9

f(x)=\frac{1}{6}x+5

h[f(x)]=h(\frac{1}{6}x+5)

h[f(x)]=2 (\frac{1}{6}x+5)+9

h[f(6)]=2(\frac{1}{6}*6+5)+9

h[f(6)]=2(1+5)+9

h[f(6)]=12+9

h[f(6)]=21